OpenStudy (anonymous):

Solve to have a general Solutions and then apply Boundary Please this time someone who can explain

OpenStudy (anonymous):

\[\frac{ d^2y }{ dx^2 }+y=0\] Boundaries are \[y(0)=1\] And \[y'(\frac{ \pi }{ 2 })= -1\]

OpenStudy (phi):

what technique were you taught to solve this type problem ?

OpenStudy (anonymous):

Separation variable method for general solution

OpenStudy (anonymous):

Coefficient , comparison and substitute method for boundaries

OpenStudy (phi):

how far did you get ?

OpenStudy (phi):

ok, but I would flip both sides

OpenStudy (anonymous):